OpenStudy (anonymous):

we have 3 root 75 minus root 243?

OpenStudy (anonymous):

\[\sqrt[3]{75} - \sqrt{243} , is this the question?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or \[3 \sqrt{75} - \sqrt{243}\] ? which one ?

OpenStudy (anonymous):

\[3\sqrt{75} - \sqrt{243} = 3\sqrt{25}\sqrt{3} - \sqrt{81}\sqrt{3} = 15\sqrt{3}-9\sqrt{3} = 6\sqrt{3}\]